OpenStudy (savannah_noelle):

Can anyone help me with 3 algebra problems please?

OpenStudy (anonymous):

Sure

OpenStudy (savannah_noelle):

thank you :)

OpenStudy (anonymous):

woah sorry i thought that was algebra 2 sry can't help but @thomaster can

OpenStudy (savannah_noelle):

Awwuh okay, well thanks for trying :)

OpenStudy (anonymous):

ur welcome good luck

OpenStudy (texaschic101):

y = -3x + 8....the slope of this line is -3. However, we need a perpendicular line, so we need the negative reciprocal of that slope. All that means is " flip " the slope and change the sign. slope -3, or -3/1....flip it and change the sign, making it 1/3. y = mx + b slope(m) = 1/3 (3,4) x = 3 and y = 4 now sub 4 = 1/3(3) + b 4 = 1 + b 4 - 1 = b 3 = b your perpendicular line is : y = 1/3x + 3 Now if you were looking for a parallel line, the slope would stay the same as in the original equation...it would be 3.

OpenStudy (texaschic101):

oops...if it were parallel, the slope would stay the same and be -3

OpenStudy (texaschic101):

any questions ?

OpenStudy (savannah_noelle):

Every time they ask for the perpendicular line that just means to use the reciprocal?

OpenStudy (texaschic101):

use the negative reciprocal....flip the slope and change the sign.. slope -4.....negative reciprocal is 1/4 slope 4/3....negative reciprocal is -3/4 slope -1/5...negative reciprocal is 5 understand ?

OpenStudy (texaschic101):

do not forget to change the sign...it makes a difference

OpenStudy (texaschic101):

you got more ?

OpenStudy (savannah_noelle):

yes I do, I tried to do this myself with the way you taught me above but it says its wrong

OpenStudy (texaschic101):

y = 4x - 1....a parallel line will have the same slope. Do you know the slope of this equation ?

OpenStudy (anonymous):

Two fair dice are rolled, and the product of the two numbers is recorded. Which of the following best describes the theoretical probability distribution?

OpenStudy (savannah_noelle):

The slope is 4 @texaschic101

OpenStudy (texaschic101):

y = mx + b slope(m) = 4 (2,0) x = 2 and y = 0 now sub 0 = 4(2) + b 0 = 8 + b -8 = b equation is : y = 4x - 8

OpenStudy (savannah_noelle):

Thank you, I was putting y=4x-4 which is why I got it wrong :/ there's one more

OpenStudy (savannah_noelle):

@taffalousswagg_bru

OpenStudy (texaschic101):

what you have is a horizontal line. A horizontal line is written as y = ...... so the equation for this one is y = 1......y is never going to be different.

OpenStudy (savannah_noelle):

@texaschic101 Thank you! :)

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

hold on im working the problem out